OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of thirty-seven and five-sevenths inches. What is its diameter? Use twenty-two sevenths for pi. twelve inches eighteen inches twenty-one inches twenty-six inches

OpenStudy (accessdenied):

We would need the formula for Circumference of a circle in terms of its diameter. Do you have that equation?

OpenStudy (anonymous):

no

OpenStudy (accessdenied):

\( C = 2 \pi r = \pi d \) Does this equation seem familiar then?

OpenStudy (anonymous):

no not really srry

OpenStudy (accessdenied):

It is fine. I recommend knowing this one, though, because it is important for these types of problems. With that, we had some information from the word problem that we can substitute in to find d. The circle formed by the blades of a fan has a[[ circumference of thirty-seven and five-sevenths inches]]. What is its diameter? Use [[twenty-two sevenths for pi.]]

OpenStudy (accessdenied):

What does "circumference is thirty-seven and five-sevenths" translate to with symbols? (Like, C = ..)

OpenStudy (anonymous):

wat? sry didnt understand that

OpenStudy (accessdenied):

Our goal is to solve for d from this equation: \( C = \pi d \) From this information, we can find a value for C: "circumference is thirty-seven and five-sevenths" That value of C goes into the equation: \( \_\_ = \pi d \)

OpenStudy (anonymous):

oh ok then wat

OpenStudy (accessdenied):

then we use the approximation for \( \pi \): \( \pi \approx \dfrac{22}{7} \) \( \_\_ = \dfrac{22}{7} d \) where the blank is filled out with the value of circumference. We simply solve for d by removing its coefficient.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then wat

OpenStudy (accessdenied):

you would be done. The value of d will be the diameter, the answer.

OpenStudy (anonymous):

but wat is the the value of d?

OpenStudy (accessdenied):

Did you follow the steps above? \( C = \pi d \) \( C = \dfrac{22}{7} d \) <-- substitute C=__ here, then solve for d . \( \_\_ = \dfrac{22}{7} d \)

OpenStudy (anonymous):

i dont get it

OpenStudy (accessdenied):

which part does not make sense?

OpenStudy (anonymous):

how am i gonig to solve for d?

OpenStudy (anonymous):

i dont get the way u put it

OpenStudy (accessdenied):

let me back up, what is the value of our circumference given in the problem? this is what goes in the blank.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then

OpenStudy (accessdenied):

you haven't answered the question: what is the value of our circumference given in the problem?

OpenStudy (anonymous):

its 37, 5/7

OpenStudy (accessdenied):

yup. so we put this in for the blank 37 5/7 = 22/7 d now we have to divide both sides by the 22/7 that is multiplied to d.

OpenStudy (accessdenied):

37 5/7 / (22/7) = (22/7) / (22/7) d divided 22/7 from both sides 37 5/7 / (22/7) = d that makes sense?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

where do i divided 22/7?

OpenStudy (anonymous):

i just see a bunch of numbers

OpenStudy (accessdenied):

let me go back and try to make this easier: \( 37 \dfrac{5}{7} = \dfrac{22}{7} d \) we should convert 37 5/7 to a mixed fraction. we'd multiply 37 by 7 and add it to the numerator like this: \( \dfrac{37\times 7 + 5}{7} = \dfrac{22}{7} d \) \( \dfrac{264}{7} = \dfrac{22}{7} d \) it might be easier, rather than do that full step, to just take it in smaller steps: \( \dfrac{264}{\cancel{7}} \color{green}{\times \cancel{7}} = \dfrac{22}{\cancel{7}} d \color{green}{\times \cancel{7}} \) \( 264 = 22 d \)

OpenStudy (anonymous):

ok then wat?

OpenStudy (accessdenied):

the last step is dividing both sides by the 22, thus getting d alone. see how to do that?

OpenStudy (accessdenied):

\( \dfrac{264}{22} = \dfrac{\cancel{22}}{\cancel{22}} d \) this is something for a calculator: 264/22, but you should find it comes out nicely

OpenStudy (anonymous):

oh ok thx so much 4 ur help

OpenStudy (anonymous):

the answer is 12in.

OpenStudy (accessdenied):

yup. glad to help. :)